# Packing subgraphs in regular graphs

### v1

@@ 3.3 The regularity lemma and the blow-up lemma
The sets $V_{1},\ldots,V_{m}$ in Lemma˜3.5 are called clusters, and the set $V_{0}$ is called the exceptional set. Given clusters and $G^{\prime}$ as in Lemma 3.5, the reduced graph $\Gamma$ is the graph whose vertices are $V_{1},\ldots,V_{m}$ (note that $V_{0}$ is omitted here) and in which $V_{i}$ is adjacent to $V_{j}$ whenever $G^{\prime}[V_{i},V_{j}]$ is $\varepsilon$-regular and has density more than $\mu$. Thus, $V_{i}V_{j}$ is an edge of $\Gamma$ if and only if $G^{\prime}$ has an edge between $V_{i}$ and $V_{j}$.
@@ 4 Balancing the expanders
As discussed in the proof sketch, our proof of Theorem˜1.4 starts by partitioning the vertex set of our dense regular graph $G$ into a small number of expanders (using Lemma 3.2). A natural strategy for finding a $K_{t,t}$-packing in $G$ is to find a $K_{t,t}$-packing in each of these expanders. However, if any of these expanders happens to be a rather imbalanced bipartite graph, then it is impossible to find a $K_{t,t}$-packing in it covering all but a constant number of vertices. To overcome this difficulty, we need the following key lemma which shows that we can partition the vertex set of any dense regular graph $G$ into a set $L$ that contains a perfect $K_{t,t}$-packing, a small set of constant size, and a small number of expanders, each of which is either far from bipartite or bipartite and balanced. Using Lemma 5.1 we can then find a $K_{t,t}$-packing in each of these expanders covering all but a constant number of its vertices, giving us the desired $K_{t,t}$-packing in $G$.
@@
Let $c,t,n_{0},n,d$ be as in the statement of the lemma, and let $G$ be a $d$-regular graph on $n$ vertices, where $d\geq cn$. By applying Lemma˜3.2 to $G$ (with $\min\{c,1/t\}$ playing the role of $c$) we obtain positive numbers $r,\eta,\beta,\gamma,\zeta,\delta$, and a partition $\{Z_{1},\ldots,Z_{r}\}$ of $V(G)$ satisfying $r\leq\left\lceil 1/c\right\rceil$ and (E1)(E4), such that
@@
Notice that this can easily be achieved by taking $\{X_{i},Y_{i}\}$ to be a random partition of $Z_{i}$. Morever, these conditions are satisfied also when $G[Z_{i}]$ is $\beta$-almost-bipartite. Indeed, if $G[Z_{i}]$ is $\beta$-almost-bipartite, then $e_{G}(X_{i})+e_{G}(Y_{i})\leq\beta n^{2}$, implying that (G1) holds. Moreover, since $\{X_{i},Y_{i}\}$ is a partition maximising the number of edges in $G[X_{i},Y_{i}]$ and since the minimum degree in $G[Z_{i}]$ is at least $\delta n$ (by (E2)), the graph $G[X_{i},Y_{i}]$ has minimum degree at least $\delta n/2$, showing (G2).
We claim that the following holds.
@@ 4.1 Procedure for moving vertices between the expanders for better degree control
Let us bound the number of steps taken by the procedure Move-high-degree-vertices by the time it ends, that is, by the time (2) is no longer satisfied by any choice of parameters. Notice that after each step of the procedure, the number of edges in $H$ decreases by at least $d-d_{H}(v)\geq\delta n/3$ (where the inequality follows because $H$ satisfies (H3) by Claim 4.2). Since $H$ satisfies (H4) at the beginning of the procedure, this implies that the total number of steps taken before the procedure ends is at most $2r\beta n^{2}/(\delta n/3)=(6r\beta/\delta)n\leq\rho n/2$ (using $r\leq\lceil 2/c\rceil$ and $\beta\ll\rho,\delta,c$). Since at most one vertex is moved at each step of the procedure we obtain the following.
@@
    At most $(6r\beta/\delta)n\leq\rho n/2$ vertices are moved during the procedure Move-high-degree-vertices.
@@
If $|\mathcal{K}^{\prime}|=\left\lfloor f/d\right\rfloor$, then we are done with the proof of the induction step, so suppose $|\mathcal{K}^{\prime}|<\left\lfloor f/d\right\rfloor$. Let $V$ be the set of all vertices contained in a copy of $K_{t,t}$ in $\mathcal{K}\cup\mathcal{K}^{\prime}$. Then $|V|\leq(2r)^{2}\cdot(f/d)\cdot 2t=8r^{2}ft/d$ because each pair $(W_{i^{\prime}},W_{j^{\prime}})$ that precedes $(W_{i},W_{j})$ in $\sigma$ satisfies $e_{H}(W_{i^{\prime}},W_{j^{\prime}})\leq f$ by the choice of $\sigma$, and there are at most $(2r)^{2}$ such pairs $(W_{i^{\prime}},W_{j^{\prime}})$. Thus, the number of edges in $H[W_{i},W_{j}]$ incident to a vertex from $V$ is at most $\rho n\cdot|V|\leq\rho n\cdot 8r^{2}ft/d\leq f/2$ (using that $\Delta(H[W_{i},W_{j}])\leq\rho n$ by (H6), that $r\leq\left\lceil 1/c\right\rceil$ and that $\rho\ll c,1/t$). This implies that there is a vertex $v$ in $W_{i}$ with at least $\frac{f}{2n}\geq\frac{Td}{2n}\geq 2t\left(\frac{4e}{\rho}\right)^{t}$ neighbours in $W_{j}-V$; denote the set of these neighbours by $A$. Observe that $G[W_{j}-V,U_{j}-V]$ has minimum degree at least $\rho n/4$ because (H5) holds (by Claim˜4.2) and $|V|\leq 2t\cdot e(H)/d\leq 2t\cdot 2r\beta n^{2}/d\leq\rho n/4$ because (H4) holds, $d\geq cn$ and $\beta\ll\rho,c,1/t$. Thus, by the Kővári–Sós–Turán theorem (Lemma˜3.8), there is a copy of $K_{t,t-1}$ in $G[A,U_{j}-V]$, with $t$ vertices in $A$ and t1t-1 vertices in $U_{j}-V$. Together with $v$, this copy of $K_{t,t-1}$ forms a copy $K$ of $K_{t,t}$ in $G$ satisfying $V(K)\subseteq Z_{i}\cup Z_{j}$ and (4), contradicting the maximality of the collection $\mathcal{K}^{\prime}$, and proving the claim. ∎
Let $\mathcal{K}$ be the collection of copies of $K_{t,t}$ guaranteed by Claim˜4.3. We claim that, for each $i\in[r]$, we have
@@
Indeed, let us fix $i\in[r]$, and note that Claim˜4.3 ensures that for every $j\in[r]$, $W_{i}\in\{X_{i},Y_{i}\}$, $W_{j}\in\{X_{j},Y_{j}\}$ such that $e_{H}(W_{i},W_{j})\geq Td$ and $W_{j}\neq Z_{i}-W_{i}$, there are exactly $\left\lfloor\frac{e_{H}(W_{i},W_{j})}{d}\right\rfloor$ copies $K\in\mathcal{K}$ satisfying $V(K)\subseteq Z_{i}\cup Z_{j}$ and (4). Therefore, it is easy to see that the following holds (where $\mathbb{1}\{E\}$ denotes the indicator function of $E$ i.e., $\mathbb{1}\{E\}=1$ if $E$ holds, and $\mathbb{1}\{E\}=0$, otherwise).
@@
$\quad=\,\sum_{W_{j}\in\{X_{j},Y_{j}\},\,\,j\in[r]-\{i\}}\left(\left\lfloor\frac{e_{H}(X_{i},W_{j})}{d}\right\rfloor\mathbb{1}\{e_{H}(X_{i},W_{j})\geq Td\}-\left\lfloor\frac{e_{H}(Y_{i},W_{j})}{d}\right\rfloor\mathbb{1}\{e_{H}(Y_{i},W_{j})\geq Td\}\right)$
$\qquad\quad+\left(2\left\lfloor\frac{e_{H}(X_{i})}{d}\right\rfloor\mathbb{1}\{e_{H}(X_{i})\geq Td\}-2\left\lfloor\frac{e_{H}(Y_{i})}{d}\right\rfloor\mathbb{1}\{e_{H}(Y_{i})\geq Td\}\right)$
@@
where for the last equality we used that (H1) holds (by Claim˜4.2). Hence (5) holds.
@@
Let us make a simple observation concerning the number of vertices added to or removed from the set $Z_{i}^{\rm{start}}$ in order to form the set $Z_{i}^{\prime}$ for every $i\in[r]$. Note that, for every $i\in[r]$, by (M) and the definition of $Z_{i}^{\prime}$, we have $|Z_{i}^{\rm{start}}-Z_{i}^{\prime}|\leq(6r\beta/\delta)n+|V(\mathcal{K})|+16rT$ and $|Z_{i}^{\prime}-Z_{i}^{\rm{start}}|\leq(6r\beta/\delta)n$. Moreover, since $r\leq\left\lceil 1/c\right\rceil$ and $\beta\ll\sigma,c,\delta$, we have $(6r\beta/\delta)n\leq\sigma n/4$, and by Claim 4.3, (H4) and the fact that $\beta\ll\sigma,c,1/t$, we have $|V(\mathcal{K})|\leq 2t\cdot\frac{e(H)}{d}\leq 2t\cdot\frac{2r\beta n^{2}}{d}\leq\sigma n/4$. Therefore,
@@ 4.3 Showing that expansion is preserved after balancing
Indeed, Claim˜4.4 shows (a).
For proving (b), notice that if $G[Z_{i}^{\textrm{start}}]$ is $\beta$-almost-bipartite, then $G^{\prime}[Z_{i}^{\prime}]$ is indeed bipartite and balanced, by the choice of $X_{i}^{\prime}$ and $Y_{i}^{\prime}$. So consider $i\in[r]$ such that $G[Z_{i}^{\textrm{start}}]$ is $\gamma$-far-from-bipartite. Recall that $|Z_{i}^{\textrm{start}}-Z_{i}^{\prime}|\leq\sigma n\leq\gamma n/4$ by (7) and the fact that $\sigma\ll\gamma$. Suppose for a contradiction that one can remove fewer than $\gamma n^{2}/4$ edges from $G^{\prime}[Z_{i}^{\prime}]=G[Z_{i}^{\prime}]$ to make it bipartite. But this means one can remove fewer than $\gamma n^{2}/4+|Z_{i}^{\textrm{start}}-Z_{i}^{\prime}|\cdot n\leq\gamma n^{2}/2$ edges from $G[Z_{i}^{\textrm{start}}]$ to make it bipartite. (Indeed, we can remove all edges incident to $Z_{i}^{\textrm{start}}-Z_{i}^{\prime}$, and fewer than $\gamma n^{2}/4$ edges from $G[Z_{i}^{\textrm{start}}\cap Z_{i}^{\prime}]$ to make it bipartite.) This contradicts $G[Z_{i}^{\textrm{start}}]$ being $\gamma$-far-from-bipartite, and proves (b).
@@
where in the last inequality we used $|Z_{i}^{\prime}|\geq\rho n/4$, which follows from $G[Z_{i}^{\prime}]$ having minimum degree at least $\rho n/4$, as pointed out at the beginning of the proof of Claim˜4.4. Hence, by (10), the average degree of $G^{\prime}[Z_{i}^{\prime}]$ is at least $d-\frac{144\sigma}{\rho^{2}}|Z_{i}^{\prime}|$, proving (c).
@@ 5 Packing $K_{t,t}$’s in expanders
Note that the following corollary, where we drop the assumption of divisibility from Lemma˜5.1 and allow for up to 2t12t-1 uncovered vertices, follows directly from Lemma˜5.1. Indeed, given $G$ as in the next corollary, remove up to 2t12t-1 vertices to obtain a graph $G^{\prime}$ whose number of vertices $n^{\prime}$ is divisible by $2t$. It is easy to see that the properties (Q1)(Q3) above hold for $G^{\prime}$, with slightly worse parameters, say, with $n_{0}-2t,c/2,\gamma/2,2\eta,\zeta/2,n^{\prime},d$ in place of $n_{0},c,\gamma,\eta,\zeta,n,d$ respectively.
@@
Let us now proceed with the proof of Lemma˜5.1.
@@
Apply the regularity lemma (Lemma˜3.5) to $G$ with parameters $\varepsilon$ and $\mu$, to obtain positive numbers $M=M(\varepsilon)$, $m$ with $m\leq M$, a subgraph $G^{\prime}$ and a partition $\{V_{0},\ldots,V_{m}\}$ of $V(G)$ into clusters $V_{1},\ldots,V_{m}$ and an exceptional set $V_{0}$ satisfying properties (R1)(R6). Let $\Gamma$ be the corresponding reduced graph, namely, it has vertex set $\{V_{1},\ldots,V_{m}\}$ and $V_{i}V_{j}$ is an edge in $\Gamma$ if and only if $G^{\prime}[V_{i},V_{j}]$ is $\varepsilon$-regular with density more than $\mu$. Equivalently, $V_{i}V_{j}$ is an edge in $\Gamma$ if and only if there is an edge of $G^{\prime}$ between $V_{i}$ and $V_{j}$ (by (R6)).
@@
In the rest of the proof, we will assume that $m$ is even, and that if $G$ is bipartite and balanced, then both $G^{\prime}$ and $\Gamma$ are bipartite and balanced. Indeed, since all clusters $V_{i}$ with $i\in[m]$ are equal in size — each of size at most $\varepsilon n$ by (R3) — and each $V_{i}$ lies entirely within one of the two parts of the bipartition of $G$ by (R2), this adjustment can be made, if necessary, by moving the vertices of a single cluster $V_{i}$ (for some $i\in[m]$) from $G^{\prime}$ into $V_{0}$. Moreover, after (potentially) moving these vertices, it still follows from (R3) that
@@
using $|X|\leq n/2$ for the third inequality, and $\varepsilon,\mu\ll\zeta$ for the last inequality. This proves (G’2). Finally, to prove (G’3), observe that if $G$ is bipartite and balanced, then, as noted before, we can assume that $G^{\prime}$ is also bipartite and balanced. Otherwise, by our assumption in Lemma˜5.1, $G$ is $\gamma$-far-from-bipartite, in which case we claim that at least $\gamma n^{2}/2$ edges need to be removed from $G^{\prime}$ to make it bipartite. Indeed, by (11) and (12), the number of edges in $G$ that are not present in $G^{\prime}$ is at most $|V_{0}|\cdot n+n\cdot(\mu+2\varepsilon)n\leq(\mu+4\varepsilon)n^{2}\leq\gamma n^{2}/2$. Since $G$ is $\gamma$-far-from-bipartite, at least $\gamma n^{2}$ edges must be removed to make it bipartite. Therefore, even after accounting for the fact that at most $\gamma n^{2}/2$ of these edges are missing in $G^{\prime}$, one still needs to remove at least $\gamma n^{2}/2$ edges from $G^{\prime}$, proving (G’3).

### v2

@@ 3.3 The regularity lemma and the blow-up lemma
The sets $V_{1},\ldots,V_{m}$ in Lemma 3.5 are called clusters, and the set $V_{0}$ is called the exceptional set. Given clusters and $G^{\prime}$ as in Lemma 3.5, the reduced graph $\Gamma$ is the graph whose vertices are $V_{1},\ldots,V_{m}$ (note that $V_{0}$ is omitted here) and in which $V_{i}$ is adjacent to $V_{j}$ whenever $G^{\prime}[V_{i},V_{j}]$ is $\varepsilon$-regular and has density more than $\mu$. Thus, $V_{i}V_{j}$ is an edge of $\Gamma$ if and only if $G^{\prime}$ has an edge between $V_{i}$ and $V_{j}$.
@@ 4 Balancing the expanders
As discussed in the proof sketch, our proof of Theorem 1.4 starts by partitioning the vertex set of our dense regular graph $G$ into a small number of expanders (using Lemma 3.2). A natural strategy for finding a $K_{t,t}$-packing in $G$ is to find a $K_{t,t}$-packing in each of these expanders. However, if any of these expanders happens to be a rather imbalanced bipartite graph, then it is impossible to find a $K_{t,t}$-packing in it covering all but a constant number of vertices. To overcome this difficulty, we need the following key lemma which shows that we can partition the vertex set of any dense regular graph $G$ into a set $L$ that contains a perfect $K_{t,t}$-packing, a small set of constant size, and a small number of expanders, each of which is either far from bipartite or bipartite and balanced. Using Corollary 5.2 we can then find a $K_{t,t}$-packing in each of these expanders covering all but a constant number of its vertices, giving us the desired $K_{t,t}$-packing in $G$.
@@
Let $c,t,n_{0},n,d$ be as in the statement of the lemma, and let $G$ be a $d$-regular graph on $n$ vertices, where $d\geq cn$. By applying Lemma 3.2 to $G$ (with $\min\{c,1/t\}$ playing the role of $c$) we obtain positive numbers $r,\eta,\beta,\gamma,\zeta,\delta$, and a partition $\{Z_{1},\ldots,Z_{r}\}$ of $V(G)$ satisfying $r\leq\left\lceil 1/c\right\rceil$ and (E1)(E4), such that
@@
Notice that this can easily be achieved by taking $\{X_{i},Y_{i}\}$ to be a random partition of $Z_{i}$. Moreover, these conditions are satisfied also when $G[Z_{i}]$ is $\beta$-almost-bipartite. Indeed, if $G[Z_{i}]$ is $\beta$-almost-bipartite, then $e_{G}(X_{i})+e_{G}(Y_{i})\leq\beta n^{2}$, implying that (G1) holds. Moreover, since $\{X_{i},Y_{i}\}$ is a partition maximising the number of edges in $G[X_{i},Y_{i}]$ and since the minimum degree in $G[Z_{i}]$ is at least $\delta n$ (by (E2)), the graph $G[X_{i},Y_{i}]$ has minimum degree at least $\delta n/2$, showing (G2).
As mentioned in the proof sketch, we would like to use the regularity of $G$ to correct the imbalance between $X_{i}$ and $Y_{i}$, for every $i\in[r]$. This will be done via the following equation, which we claim holds for every $i\in[r]$, and which generalises (1) from the proof sketch from bipartite graphs to general graphs.
@@
As explained in the proof sketch (see also Claim 4.3 below), (2) shows that to find a balancing $K_{t,t}$-packing, it suffices to find disjoint $K_{t,t}$-packings for each pair $(W_{i},W_{j})$ with $W_{i}\in\{X_{i},Y_{i}\}$ and $W_{j}\in\{X_{j},Y_{j}\}$ of size roughly $\frac{1}{d}e_{G}(W_{i},W_{j})$, whose $K_{t,t}$’s effectively move one vertex from $W_{i}$ to $Z_{j}-W_{j}$. This can be done greedily if the maximum degree of edges between expanders is small (at most $\rho n$ for a small constant $\rho$). However, the maximum degree of edges between expanders need not be small, even if we are allowed to modify the $Z_{i}$’s somewhat. Nevertheless, we emulate the small maximum degree situation by constructing a subgraph $H$ of $G$, and modifying the sets $Z_{i}$, so that (2) is maintained and the maximum degree of the final $H$ is small (with possibly some, very few, exceptional vertices with large degree). To do so we first define $H$ as follows, and then modify it and the $Z_{i}$’s as described in the next subsection.
@@ 4.1 Procedure for moving vertices between the expanders for better degree control
Let us bound the number of steps taken by the procedure Move-high-degree-vertices by the time it ends, that is, by the time (3) is no longer satisfied by any choice of parameters. Notice that after each step of the procedure, the number of edges in $H$ decreases by at least $\rho n$. Since $H$ satisfies (H4) at the beginning of the procedure, this implies that the total number of steps taken before the procedure ends is at most $2r\beta n^{2}/(\rho n)=(2r\beta/\rho)n\leq\rho n/2$ (using $r\leq\lceil 2/c\rceil$ and $\beta\ll\rho,\delta,c$). Since at most one vertex is moved at each step of the procedure we obtain the following.
@@
    At most $(2r\beta/\rho)n\leq\rho n/2$ vertices are moved during the procedure Move-high-degree-vertices.
@@
If $|\mathcal{K}^{\prime}|=\left\lfloor f/d\right\rfloor$, then we are done with the proof of the induction step, so suppose $|\mathcal{K}^{\prime}|<\left\lfloor f/d\right\rfloor$. Let $V$ be the set of all vertices contained in a copy of $K_{t,t}$ in $\mathcal{K}\cup\mathcal{K}^{\prime}$. Then $|V|\leq(2r)^{2}\cdot(f/d)\cdot 2t=8r^{2}ft/d$ because each pair $(W_{i^{\prime}},W_{j^{\prime}})$ that precedes $(W_{i},W_{j})$ in $\sigma$ satisfies $e_{H}(W_{i^{\prime}},W_{j^{\prime}})\leq f$ by the choice of $\sigma$, and there are at most $(2r)^{2}$ such pairs $(W_{i^{\prime}},W_{j^{\prime}})$. Thus, the number of edges in $H[W_{i},W_{j}]$ incident to a vertex from $V$ is at most $\rho n\cdot|V|\leq\rho n\cdot 8r^{2}ft/d\leq f/2$ (using that $\Delta(H[W_{i},W_{j}])\leq\rho n$ by (H6), that $r\leq\left\lceil 1/c\right\rceil$ and that $\rho\ll c,1/t$). This implies that there is a vertex $v$ in $W_{i}$ with at least $\frac{f}{2n}\geq\frac{Td}{2n}\geq 2t\left(\frac{4e}{\rho}\right)^{t}$ neighbours in $W_{j}-V$; denote the set of these neighbours by $A$. Observe that $G[W_{j}-V,U_{j}-V]$ has minimum degree at least $\rho n/4$ because (H5) holds (by Claim 4.2) and $|V|\leq 2t\cdot e(H)/d\leq 2t\cdot 2r\beta n^{2}/d\leq\rho n/4$ because (H4) holds, $d\geq cn$ and $\beta\ll\rho,c,1/t$. Thus, by the Kővári–Sós–Turán theorem (Lemma 3.8), there is a copy of $K_{t,t-1}$ in $G[A,U_{j}-V]$, with $t$ vertices in $A$ and t1t-1 vertices in $U_{j}-V$. Together with $v$, this copy of $K_{t,t-1}$ forms a copy $K$ of $K_{t,t}$ in $G$ satisfying $V(K)\subseteq Z_{i}\cup Z_{j}$ and (5), contradicting the maximality of the collection $\mathcal{K}^{\prime}$, and proving the claim. ∎
Let $\mathcal{K}$ be the collection of copies of $K_{t,t}$ guaranteed by Claim 4.3. We claim that, for each $i\in[r]$, we have
@@
Indeed, let us fix $i\in[r]$, and note that Claim 4.3 ensures that for every $j\in[r]$, $W_{i}\in\{X_{i},Y_{i}\}$, $W_{j}\in\{X_{j},Y_{j}\}$ such that $e_{H}(W_{i},W_{j})\geq Td$ and $W_{j}\neq Z_{i}-W_{i}$, there are exactly $\left\lfloor\frac{e_{H}(W_{i},W_{j})}{d}\right\rfloor$ copies $K\in\mathcal{K}$ satisfying $V(K)\subseteq Z_{i}\cup Z_{j}$ and (5). Therefore, it is easy to see that the following holds (where $\mathbbm{1}\{E\}$ denotes the indicator function of $E$ i.e., $\mathbbm{1}\{E\}=1$ if $E$ holds, and $\mathbbm{1}\{E\}=0$, otherwise).
@@
$\quad=\,\sum_{W_{j}\in\{X_{j},Y_{j}\},\,\,j\in[r]-\{i\}}\left(\left\lfloor\frac{e_{H}(X_{i},W_{j})}{d}\right\rfloor\mathbbm{1}\{e_{H}(X_{i},W_{j})\geq Td\}-\left\lfloor\frac{e_{H}(Y_{i},W_{j})}{d}\right\rfloor\mathbbm{1}\{e_{H}(Y_{i},W_{j})\geq Td\}\right)$
$\qquad\quad+\left(2\left\lfloor\frac{e_{H}(X_{i})}{d}\right\rfloor\mathbbm{1}\{e_{H}(X_{i})\geq Td\}-2\left\lfloor\frac{e_{H}(Y_{i})}{d}\right\rfloor\mathbbm{1}\{e_{H}(Y_{i})\geq Td\}\right)$
@@
where for the last equality we used that (H1) holds (by Claim 4.2). Hence (6) holds.
@@
Let us make a simple observation concerning the number of vertices added to or removed from the set $Z_{i}^{\rm{start}}$ in order to form the set $Z_{i}^{\prime}$ for every $i\in[r]$. Note that, for every $i\in[r]$, by (M) and the definition of $Z_{i}^{\prime}$, we have $|Z_{i}^{\rm{start}}-Z_{i}^{\prime}|\leq(2r\beta/\rho)n+|V(\mathcal{K})|+16rT$ and $|Z_{i}^{\prime}-Z_{i}^{\rm{start}}|\leq(2r\beta/\rho)n$. Moreover, since $r\leq\left\lceil 1/c\right\rceil$ and $\beta\ll\rho,\sigma,c,\delta$, we have $(2r\beta/\rho)n\leq\sigma n/4$, and by Claim 4.3, (H4) and the fact that $\beta\ll\rho,\sigma,c,1/t$, we have $|V(\mathcal{K})|\leq 2t\cdot\frac{e(H)}{d}\leq 2t\cdot\frac{2r\beta n^{2}}{d}\leq\sigma n/4$. Therefore,
@@ 4.3 Showing that expansion is preserved after balancing
Indeed, Claim 4.4 shows (a).
For proving (b), notice that if $G[Z_{i}^{\textrm{start}}]$ is $\beta$-almost-bipartite, then $G^{\prime}[Z_{i}^{\prime}]$ is indeed bipartite and balanced, by the choice of $X_{i}^{\prime}$ and $Y_{i}^{\prime}$. So consider $i\in[r]$ such that $G[Z_{i}^{\textrm{start}}]$ is $\gamma$-far-from-bipartite. Recall that $|Z_{i}^{\textrm{start}}-Z_{i}^{\prime}|\leq\sigma n\leq\gamma n/4$ by (8) and the fact that $\sigma\ll\gamma$. Suppose for a contradiction that one can remove fewer than $\gamma n^{2}/4$ edges from $G^{\prime}[Z_{i}^{\prime}]=G[Z_{i}^{\prime}]$ to make it bipartite. But this means one can remove fewer than $\gamma n^{2}/4+|Z_{i}^{\textrm{start}}-Z_{i}^{\prime}|\cdot n\leq\gamma n^{2}/2$ edges from $G[Z_{i}^{\textrm{start}}]$ to make it bipartite. (Indeed, we can remove all edges incident to $Z_{i}^{\textrm{start}}-Z_{i}^{\prime}$, and fewer than $\gamma n^{2}/4$ edges from $G[Z_{i}^{\textrm{start}}\cap Z_{i}^{\prime}]$ to make it bipartite.) This contradicts the fact that $G[Z_{i}^{\textrm{start}}]$ is $\gamma$-far-from-bipartite, and proves (b).
@@
where in the last inequality we used $|Z_{i}^{\prime}|\geq\rho n/4$, which follows from $G[Z_{i}^{\prime}]$ having minimum degree at least $\rho n/4$, as pointed out at the beginning of the proof of Claim 4.4. Hence, by (11), the average degree of $G^{\prime}[Z_{i}^{\prime}]$ is at least $d-\frac{144\sigma}{\rho^{2}}|Z_{i}^{\prime}|$, proving (c).
@@ 5 Packing $K_{t,t}$’s in expanders
Note that the following corollary, where we drop the assumption of divisibility from Lemma 5.1 and allow for up to 2t12t-1 uncovered vertices, follows directly from Lemma 5.1. Indeed, given $G$ as in the next corollary, remove up to 2t12t-1 vertices to obtain a graph $G^{\prime}$ whose number of vertices $n^{\prime}$ is divisible by $2t$. It is easy to see that the properties (Q1)(Q3) above hold for $G^{\prime}$, with slightly worse parameters, say, with $n_{0}-2t,c/2,\gamma/2,2\eta,\zeta/2,n^{\prime},d$ in place of $n_{0},c,\gamma,\eta,\zeta,n,d$ respectively.
@@
Let us now proceed with the proof of Lemma 5.1.
@@
Apply the regularity lemma (Lemma 3.5) to $G$ with parameters $\varepsilon$ and $\mu$, to obtain positive numbers $M=M(\varepsilon)$, $m$ with $m\leq M$, a subgraph $G^{\prime}$ and a partition $\{V_{0},\ldots,V_{m}\}$ of $V(G)$ into clusters $V_{1},\ldots,V_{m}$ and an exceptional set $V_{0}$ satisfying properties (R1)(R6). Let $\Gamma$ be the corresponding reduced graph, namely, it has vertex set $\{V_{1},\ldots,V_{m}\}$ and $V_{i}V_{j}$ is an edge in $\Gamma$ if and only if $G^{\prime}[V_{i},V_{j}]$ is $\varepsilon$-regular with density more than $\mu$. Equivalently, $V_{i}V_{j}$ is an edge in $\Gamma$ if and only if there is an edge of $G^{\prime}$ between $V_{i}$ and $V_{j}$ (by (R6)).
@@
In the rest of the proof, we will assume that $m$ is even, and that if $G$ is bipartite and balanced, then both $G^{\prime}$ and $\Gamma$ are bipartite and balanced. Indeed, since all clusters $V_{i}$ with $i\in[m]$ are equal in size — each of size at most $\varepsilon n$ by (R3) — and each $V_{i}$ lies entirely within one of the two parts of the bipartition of $G$ by (R2), this adjustment can be made, if necessary, by moving the vertices of some clusters, covering up to $\varepsilon n$ vertices, from $G^{\prime}$ into $V_{0}$. Moreover, after (potentially) moving these vertices, it still follows from (R3) that
@@
using $|X|\leq n/2$ for the third inequality, and $\varepsilon,\mu\ll\zeta$ for the last inequality. This proves (G’2). Finally, to prove (G’3), observe that if $G$ is bipartite and balanced, then, as noted before, we can assume that $G^{\prime}$ is also bipartite and balanced. Otherwise, by our assumption in Lemma 5.1, $G$ is $\gamma$-far-from-bipartite, in which case we claim that at least $\gamma n^{2}/2$ edges need to be removed from $G^{\prime}$ to make it bipartite. Indeed, by (12) and (13), the number of edges in $G$ that are not present in $G^{\prime}$ is at most $|V_{0}|\cdot n+n\cdot(\mu+2\varepsilon)n\leq(\mu+4\varepsilon)n^{2}\leq\gamma n^{2}/2$. Since $G$ is $\gamma$-far-from-bipartite, at least $\gamma n^{2}$ edges must be removed to make it bipartite. Therefore, even after accounting for the fact that at most $\gamma n^{2}/2$ of these edges are missing in $G^{\prime}$, one still needs to remove at least $\gamma n^{2}/2$ edges from $G^{\prime}$, proving (G’3).